# How to Evaluate Distributed Coordination Systems? – A Survey and Analysis

### v2

@@ 4.2 Availability
UPB (Under Pressure Benchmark) [29] is the benchmark for measuring and quantifying the availability of distributed database systems. UPB uses distributed YCSB workloads and a more complex evaluation scenario with different load sizes, replication factors, and the number of failed nodes. UPB aims to measure the performance impact on the system before, during, and after node failures. UPB leverages the YCSB workload generator for its availability evaluation and therefore inherits all its limitations too. That being said, UPB can be a guide for any benchmarking suit on quantifying and comparing availability.
Chaos Monkey [66] is a tool that randomly terminates virtual machine instances and containers that run inside a cloud. Such random node termination could potentially reveal any problem in distributed systems thus ensuring that engineers implement their services to be resilient to instance failures. Chaos Monkey is a good example that follows chaos engineering for fault injection in any benchmarking suit. Chaos Monkey lets the user define their own outage checker for availability checking instead of giving any performance or other evaluation as plugins.
@@ 4.3 Consistency
The benchmarking frameworks that try to evaluate consistency, like YCSB++ [70], can usually measure only the read operation staleness in a weakly consistent system. YCSB++ coordinates multiple distributed clients using ZooKeeper and measures time-based staleness. The write operation is published to ZooKeeper right after it is completed. Then any subscribing client can get the written value and repeat read operations until it reads a new value. The time difference between the first and last read is the approximated lower bound of staleness. One benefit of such a method is the ability of online measuring of consistency, and the longer the benchmark runs, the better the chance to minimize the coordination noise. Similarly, BG measures the unpredictable data amount as it constrains the synchronization of the data with an unpredictable time amount in the SLA. To measure this, BG logs each read and write request with unique IDs and then it compares their values based on the timestamp of the operation. This approach brings the accuracy for the time synchronization in the distributed setting of benchmarking clients. BenchFoundry follows a similar approach by logging all fine-grained results to evaluate staleness. On the other hand, it relies on the clock synchronization of the clients which is challenging in distributed settings. Out of the listed benchmarking tools in table 2.2, only Jepsen [47] and Elle [48] provide testing for consistency models in terms of linearizability and serializability. Jepsen does not provide a black box benchmarking interface, and it requires customizing workload characteristics depending on the consistency models. It simply enables the user to customize the workload to check whether the proposed consistency guarantees are preserved. Elle, however, can evaluate isolation guarantees of a black box distributed database more efficiently than Jepsen’s KNOSSOS [46] by constructing a dependency graph from transaction history and identifying violating cycles.
@@ 5 Related Work
Distributed systems have been maintaining their importance for the last several decades due to the increase in the need for scalable and reliable distributed applications while preserving high performance. To analyze distributed systems comprehensively and compare them in terms of features and services, various surveys and evaluations published in the past. Surveys on cloud providers, data warehouses, distributed file systems, or metadata services can be counted among them.
Cloud providers are analyzed and evaluated in terms of elasticity [79], computing power [64], and cost to performance efficiency [34] in previous efforts. Widely used distributed services are also analyzed in many works such as a survey on stream processing [75] or performance and dependability evaluation of MapReduce systems [74]. Similarly, different aspects of distributed systems are studied in several surveys like reliability analysis on distributed systems [10] and load balancing characteristics of known systems [43].
@@
Analysis of distributed coordination services in terms of general characteristics and importance of coordination [69] and the comparison of existing algorithms [51] are among the published works. However, as of our knowledge, there is no published work on the evaluation of distributed coordination systems. As mentioned in the Introduction, due to the lack of standard benchmarking tools for distributed coordination services, developers widely use their ad-hoc benchmarks which are prone to unfair comparisons or limited results for the evaluation of the systems. This study is unique in identifying the metrics and parameters for the evaluation of distributed coordination systems, discussing how each system uses these metrics and parameters for their evaluation, pinpointing the deficiencies of well-known benchmarking suites in evaluating distributed computing systems, and finally discussing the features of an ideal distributed coordination benchmark.
@@ 6 Conclusion
Increasing demand in web-based big-data applications brought out the need for efficient use of distributed systems, which highly depend on the adequate implementation of distributed coordination. The distributed systems community has developed different protocols, coordination services, and distributed applications built on top of these services to satisfy this rapid growth in big data applications. However, due to the lack of a standard benchmarking tool, developers generally opt to use ad-hoc evaluation mechanisms and microbenchmarks. Hence, the evaluation of these systems has been very limited, resulting in inadequate and misleading measurements and an unfair comparison of the competing systems. In this paper, we have analyzed and compared well-known and widely used distributed coordination services, their evaluation mechanisms, and the tools used to benchmark those systems. We have identified the essential requirements of distributed coordination service benchmarking, such as the metrics and parameters for the evaluation of the performance, scalability, availability, and consistency of these systems
Despite the commonalities in tuning parameters like read/write ratio, data access overlap, or size of data objects for performance measurements, access locality and number of data objects are not extensively evaluated by most of the systems. These two parameters, on the other hand, may have a significant impact on the system performance for systems that allow concurrent writes since they are the main factors in producing conflicting or colliding commands. Similarly, they are critical for managing ownership of data objects for systems using ownership mechanisms to resolve conflicts. While most of the systems are evaluated with a single value of read/write ratio and size of data objects, tuning these parameters for different values is essential to understand the system performance extensively for different types of workloads. Depending on the type of applications using these services, the percentage of update operations or the size of data objects may change dramatically. Likewise, data access overlap is evaluated only for systems sharing the entire data space, but it is also a system-specific parameter that needs to be analyzed for different scenarios.

### v3

@@ 4.2 Availability
UPB (Under Pressure Benchmark) [31] is the benchmark for measuring and quantifying the availability of distributed database systems. UPB uses distributed YCSB workloads and a more complex evaluation scenario with different load sizes, replication factors, and the number of failed nodes. UPB aims to measure the performance impact on the system before, during, and after node failures. UPB leverages the YCSB workload generator for its availability evaluation and therefore inherits all its limitations too. That being said, UPB can be a guide for any benchmarking suite on quantifying and comparing availability.
Chaos Monkey [68] is a tool that randomly terminates virtual machine instances and containers that run inside a cloud. Such random node termination could potentially reveal any problem in distributed systems, thus ensuring that engineers implement their services to be resilient to instance failures. Chaos Monkey is a good example that follows chaos engineering for fault injection in any benchmarking suite. Chaos Monkey lets the user define their own outage checker for availability checking instead of giving any performance or other evaluation as a plugin.
@@ 4.3 Consistency
The benchmarking frameworks that try to evaluate consistency, like YCSB++ [72], can usually measure only the read operation staleness in a weakly consistent system. YCSB++ coordinates multiple distributed clients using ZooKeeper and measures time-based staleness. The write operation is published to ZooKeeper right after it is completed. Then any subscribing client can get the written value and repeat read operations until it reads a new value. The time difference between the first and last read is the approximated lower bound of staleness. One benefit of such a method is the ability to measure consistency online, and the longer the benchmark runs, the better the chance to minimize the coordination noise. Similarly, BG measures the unpredictable data amount as it constrains the synchronization of the data with an unpredictable time amount in the SLA. To measure this, BG logs each read and write request with unique IDs, and then it compares their values based on the timestamp of the operation. This approach brings the accuracy for the time synchronization in the distributed setting of benchmarking clients. BenchFoundry follows a similar approach by logging all fine-grained results to evaluate staleness. On the other hand, it relies on the clock synchronization of the clients, which is challenging in distributed settings. Out of the listed benchmarking tools in table II, only Jepsen [49] and Elle [50] provide testing for consistency models in terms of linearizability and serializability. Jepsen does not provide a black box benchmarking interface, and it requires customizing workload characteristics depending on the consistency models. It simply enables the user to customize the workload to check whether the proposed consistency guarantees are preserved. Elle, however, can evaluate isolation guarantees of a black box distributed database more efficiently than Jepsen’s KNOSSOS [48] by constructing a dependency graph from transaction history and identifying violating cycles.
@@ 5 Related Work
Distributed systems have been maintaining their importance for the last several decades due to the increase in the need for scalable and reliable distributed applications while preserving high performance. To analyze distributed systems comprehensively and compare them in terms of features and services, various surveys and evaluations have been published in the past. Surveys on cloud providers, data warehouses, distributed file systems, or metadata services can be counted among them.
Cloud providers are analyzed and evaluated in terms of elasticity [81], computing power [66], and cost to performance efficiency [36] in previous efforts. Widely used distributed services are also analyzed in many works, such as a survey on stream processing [77] or performance and dependability evaluation of MapReduce systems [76]. Similarly, different aspects of distributed systems are studied in several surveys, like reliability analysis on distributed systems [10] and load balancing characteristics of known systems [45].
@@
Analysis of distributed coordination services in terms of general characteristics and importance of coordination [71] and the comparison of existing algorithms [53] are among the published works. However, to the best of our knowledge, there is no published work on the evaluation of distributed coordination systems. As mentioned in the Introduction, due to the lack of standard benchmarking tools for distributed coordination services, developers widely use their ad-hoc benchmarks, which are prone to unfair comparisons or limited results for the evaluation of the systems. This study is unique in identifying the metrics and parameters for the evaluation of distributed coordination systems, discussing how each system uses these metrics and parameters for its evaluation, pinpointing the deficiencies of well-known benchmarking suites in evaluating distributed computing systems, and finally discussing the features of an ideal distributed coordination benchmark.
@@ 6 Conclusion
Increasing demand in web-based big-data applications brought out the need for efficient use of distributed systems, which highly depend on the adequate implementation of distributed coordination. The distributed systems community has developed different protocols, coordination services, and distributed applications built on top of these services to satisfy this rapid growth in big data applications. However, due to the lack of a standard benchmarking tool, developers generally opt to use ad-hoc evaluation mechanisms and microbenchmarks. Hence, the evaluation of these systems has been very limited, resulting in inadequate and misleading measurements and an unfair comparison of the competing systems. In this paper, we have analyzed and compared well-known and widely used distributed coordination services, their evaluation mechanisms, and the tools used to benchmark those systems. We have identified the essential requirements of distributed coordination service benchmarking, such as the metrics and parameters for the evaluation of the performance, scalability, availability, and consistency of these systems.
Despite the commonalities in tuning parameters like read/write ratio, data access overlap, or size of data objects for performance measurements, access locality and number of data objects are not extensively evaluated by most of the systems. These two parameters, on the other hand, may have a significant impact on the system performance for systems that allow concurrent writes, since they are the main factors in producing conflicting or colliding commands. Similarly, they are critical for managing ownership of data objects for systems using ownership mechanisms to resolve conflicts. While most of the systems are evaluated with a single value of read/write ratio and size of data objects, tuning these parameters for different values is essential to understand the system performance extensively for different types of workloads. Depending on the type of applications using these services, the percentage of update operations or the size of data objects may change dramatically. Likewise, data access overlap is evaluated only for systems sharing the entire data space, but it is also a system-specific parameter that needs to be analyzed for different scenarios.
The benchmarks discussed in this work do consider some of these tuning parameters, but often do not prioritize them when evaluating coordination algorithms and services. Benchmarks on performance often neglect scalability, data access overlap, and request conflicts as seen in YCSB++ [72], BG [16], and BenchFoundry [18]. Some benchmarks on availability neglect to test for node failures, clock drift, and network partition altogether, as seen with UPB [31] and Chaos Monkey [68]. Linearizability is often challenging to evaluate and needs more attention. To overcome these limitations with the existing benchmarks, some systems (such as ScalarDB [84] and CockroachDB [5]) combines multiple existing benchmarks to achieve the desired functionality. Most of the other systems considered in this study prefer using their own specially tailored microbenchmarks instead.